# EgoPlan-Bench2: A Benchmark for Multimodal Large Language Model Planning in Real-World Scenarios

### v1

@@ I Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [1, 2, 3, 4, 5, 6, 7] has demonstrated remarkable comprehension and generalization capacities, opening new possibilities for achieving the ultimate goal of artificial general intelligence (AGI) [8, 9], which aims to match or surpass human performance in most tasks. By plugging efficient visual encoders into pretrained Large Language Models (LLMs) [10, 11, 12, 13] and learning alignments between vision and language [14], MLLMs have excelled in various multimodal tasks such as image captioning [15, 16], visual question answering [16, 17, 18, 19, 20, 21, 22], mathematical reasoning [23, 24], crossing-modality grounding [25], etc. However, achieving AGI requires more than just advanced comprehension and reasoning. A crucial milestone is attaining human-level task planning capabilities, which involve making informed decisions in complex environments. This capability is essential for developing a versatile intelligent assistant that can assist humans in tackling a wide array of real-world challenges in daily life.
@@
While the comprehension capabilities of MLLMs have been extensively evaluated in previous benchmarks [26, 27, 20, 28, 29], the evaluation of the planning abilities of current MLLMs in various scenarios remains underexplored. A comprehensive benchmark specifically designed to assess the planning capabilities of MLLMs across a wide range of real-world scenarios is highly demanded to uncover the potential of MLLMs in serving as versatile assistants in the real world. Previous egocentric video question answering (QA) benchmarks [30, 31] also evaluate model performance in everyday life, but they primarily assess comprehension rather than planning, where a model answers questions based on the spatial and temporal understanding of the entire video. Most relevant to addressing this issue is EgoPlan-Bench [32], which evaluates the planning abilities of MLLMs from an egocentric perspective. However, it is constrained to single kitchen scenarios, lacking a comprehensive evaluation across a variety of real-world contexts.
@@
In this paper, we introduce EgoPlan-Bench2, a benchmark designed to rigorously assess the planning capabilities of MLLMs across a broad range of daily scenarios. EgoPlan-Bench2 is founded on three principal tenets: a) Rich and diverse real-world scenarios. It includes 1,321 high-quality multiple-choice QA pairs sourced from 1,113 videos, covering 4 major life domains: Work, Daily life, Hobbies and Recreation. These domains are further subdivided into 24 detailed scenarios, ranging from everyday household tasks to specialized activities such as laboratory work, blacksmith and mechanical repairs as shown in Fig. 1. In addition to scenario settings, EgoPlan-Bench2 features 284 distinct verbs in task goals and 434 in candidate options, along with 742 and 1,113 unique objects respectively. The duration of task progress videos varies from a few seconds to five minutes. The rich and diverse scene setups, coupled with a variety of actions, objects and video lengths, ensure a comprehensive evaluation of MLLMs’ planning capabilities across various contexts. b) Egocentric perspective. We choose Ego4D [33] as the video source, because it provides a vast array of first-person perspective videos that capture realistic human interactions with objects and environments. This egocentric perspective is crucial for evaluating planning capabilities in a manner that closely mirrors real-world human experiences, offering a more authentic assessment of how MLLMs can assist in everyday tasks. c) Planning capability evaluation. EgoPlan-Bench2 is specifically aimed at evaluating MLLMs’ planning abilities, where a model must track long-term task progress, comprehend the current state of the environment, and leverage both general and domain-specific knowledge to plan the next action, in order to correctly answer the questions as shown in Fig. 1.
@@ II-A Advancements in Multimodal Large Language Models
Building upon the impressive achievements of LLMs, MLLMs have also experienced a revolutionary transformation. MLLMs typically consists of an image encoder (e.g., CLIP [34]) to extract visual information, a language model (e.g., LLaMA [11], Vicuna [12]) to decode multimodal or text sequence and a trainable align module (e.g., Q-Former [35], gated cross-attention layer [36]) to integrate visual features into the language embedding space. Groundbreaking models like BLIP-2 [35], LLaVA [37], Flamingo [36] and PaLM-E [38] have made early attempt to integrate LLMs into vision-language pre-training and have demonstrated remarkable multimodal understanding and reasoning capabilities. Recent research interest has increasingly shifted towards multimodal understanding and generation that incorporates videos as visual signals [39, 40, 41, 42, 43]. These methods (e.g., VideoChat [39], VideoChatGPT [40], Valley [44]) try to enhance MLLMs’ instruction-following capabilities by generating video instruction-tuning data. Video-LLaMA [45] encodes individual frames through a ViT [46] and an image Q-Former and then apply temporal process through a video Q-Former. VideoChat2 [18] encodes video frames through a video transformer, and Q-former is employed to compress video tokens. Expect these models mentioned above, a considerable number of video-based MLLMs [47, 48, 49, 50] have been proposed, demonstrating notable generalization and reasoning abilities across a wide range of tasks.
@@ II-B Benchmarking Multimodal Large Language Models
To guide the potential future development of MLLMs, numerous benchmarks have been proposed to assess model performance across various aspects and tasks. Alongside advancements in comprehensive benchmarks for image MLLMs [51, 26, 52, 53, 54, 55, 56], significant efforts have been made to creating benchmarks for video MLLMs [57, 40, 19, 18, 21]. For example, Video-MME [19] is build upon videos collected from YouTube, aiming to evaluate models’ capacities in 12 task types involving temporal perception, spatial perception, OCR, object recognition, etc. MVBench [18] focuses on temporally-sensitive videos and conducts comprehensive evaluations of MLLMs’ temporal understanding. MMWorld [21] is a benchmark characterized by its multi-discipline nature, evaluating models on multi-faceted tasks such as explanation and counterfactual thinking. MLVU [58] is a long video understanding benchmark which is built upon various video genres, including movies, cartoons, game videos, etc. The diagnostic benchmark dataset, WorldQA [59], challenges machines to answer questions about a video by employing multimodal data (auditory and visual) and world knowledge. Despite the advancement, videos in these benchmarks are not recorded from a first-person perspective and fail to simulate the realistic visual input that a model would receive during planning tasks. In addition, these benchmarks mainly focus on MLLMs’ comprehension capabilities instead of planning capabilities.
@@ II-C Egocentric Video Datasets
Numerous egocentric datasets containing daily life activity have been developed over the past years, including Ego4D [33], Epic-Kitchens [60], UT Ego [61, 62], Activities of Daily Living (ADL) [63], Disney dataset [64], Charades-Ego [65], etc. Among these egocentric datasets, we select Ego4D as the video source for EgoPlan-Bench2 due to its massive scale, encompassing hundreds of indoor and outdoor environments, a much wider demographic and an exceptionally rich variety of tasks and scene types. There are also some existing egocentric QA benchmarks like EgoThink [66], EgoVQA [30] and Egotaskqa [31], but they mainly focus on evaluating model’s reasoning and comprehension abilities of activities, human-object interactions and environments instead of task planning capacity. Most relevant to our work are EgoPlan-Bench [32] and VidEgoThink [67]. EgoPlan-Bench is also a planning benchmark that utilizes kitchen-related videos from Epic-Kitchens [60] and Ego4D datasets, but it focuses exclusively on cooking scenario and lacks evaluation across a wider variety of real task types. VidEgoThink evaluates the capabilities for different functions of MLLMs in Embodied AI from four dimensions: video QA, hierarchy planning, visual grounding and reward modeling. In the most relevant dimension of hierarchy planning, VidEgoThink includes only 9 scenes sourced from 76 videos, suffering from limited task types and scenario diversity.
@@ III-A Egocentric Video Source
Ego4D has densely timestamped annotations indicating specific actions and their occurrence time, represented by short sentences in the format of “#C C does something”. In order to remove invalid annotations and reduce noise, we adopt the criteria following the EgoVLP framework [68]:
@@
After narration filtering, we convert raw narrations to verb-object phrases like “close washing machine” with GPT-4 to better represent actions and uniform format. The raw action timestamps in Ego4D indicate when actions occur but not their durations. Following EgoVLP [68], we calculate start and end timestamps for each action. For an action with occurrence timestamp $t_{i}$, the start and end timestamps can be calculated as:
@@ III-B3 Stage III: Model and Human Verification
In the final verification phase of EgoPlan-Bench2, we exclude QA pairs that can be correctly answered using only textual information, ensuring the benchmark effectively assesses the models’ capabilities to process multimodal inputs. GPT-4 is employed with the CircularEval Strategy [26] to minimize the impact of random guessing.
@@ III-C EgoPlan-Bench2 Statistics
EgoPlan-Bench2 is highly diverse in terms of scenarios, human-object interactions, and video durations, comprising 1,321 high-quality multiple-choice QA pairs derived from 1,113 videos. Firstly, it has an extensive range of scenario settings across 4 major life domains: Work (404 QA pairs), Daily Life (439 QA pairs), Recreation (183 QA pairs) and Hobbies (295 QA pairs). A more detailed breakdown of the 24 fine-grained scenarios and the corresponding number of QA pairs is provided in Fig. 4. Second, EgoPlan-Bench2 incorporates comprehensive human-object interaction data, manifested in both actions and task goals. We use NLTK [69] to analyze the verbs and nouns in the task goals and candidate options of the QA pairs, identifying 284 distinct verbs and 742 unique objects in the task goals, and 434 verbs and 1,113 objects in the candidate options. The word clouds of these terms are displayed in Fig. 5. Third, varied video durations in EgoPlan-Bench2 range from a few seconds to five minutes. We consider 30 seconds as the threshold, classifying videos of 30 seconds or less as short videos (690 QA pairs) and those exceeding this duration as long videos (631 QA pairs).
@@ IV Experiments
Yi-VL[70] 8 Yi-6B 24.37 21.29 26.23 23.39 23.19 23.77 23.47
MultiModal-GPT[71] 8 LLaMA-7B 26.42 23.27 23.50 26.10 23.48 26.62 24.98
LLaVA1.5[72] 6 LLaMA-7B 29.61 21.04 27.32 24.07 23.77 27.26 25.44
InternVL-1.5[73, 74] 8 InternLM2-Chat-1.8B 28.02 24.75 23.50 24.41 23.91 27.42 25.59
mPLUG-Owl-2[75] 8 LLaMA2-7B 27.79 24.75 24.04 25.42 25.36 26.31 25.81
BLIP-2[35] 8 Flan-T5-XL 24.37 23.51 30.05 30.17 24.64 27.89 26.19
InstructBLIP[76] 8 Flan-T5-XL 27.33 23.02 26.23 29.49 25.51 27.26 26.34
InstructBLIP Vicuna[76] 8 Vicuna-7B 27.56 24.26 28.42 28.14 25.65 28.05 26.80
DeepSeek-VL[77] 6 DeepSeek-LLM-7B 32.12 24.75 26.23 29.83 28.55 28.53 28.54
Qwen-VL-Chat[78] 8 Qwen-7B 32.57 27.23 27.87 28.47 30.00 28.68 29.37
InternVL-2[73, 74] 8 InternLM2.5-Chat-7B 37.81 23.76 31.69 28.14 31.01 29.95 30.51
@@
Video-LLaMA2[50] 8 Mistral-v0.2-Instruct-7B 24.15 23.02 19.13 23.73 23.19 22.82 23.01
LLaVA-NeXT-Video[48] 16 Vicuna1.5-7B 26.42 19.55 24.59 23.05 22.61 24.09 23.32
Video-ChatGPT[40] 100 LLaMA-7B 24.15 22.77 24.59 24.07 23.33 24.25 23.77
Video-LLaVA[79] 8 Vicuna1.5-7B 27.11 22.52 27.87 24.75 25.22 25.36 25.28
ShareGPT4Video[47] 16 LLaMA3-Instruct-8B 25.51 23.02 26.78 27.46 25.07 25.67 25.36
VILA[80] 6 LLaMA3-8B 28.70 20.05 30.05 25.08 23.77 27.26 25.44
LongVA[49] 32 Qwen2-Instruct-7B 27.11 23.27 26.78 29.49 27.25 25.52 26.42
VideoChat2[18] 16 Mistral-v0.2-Instruct-7B 28.93 24.75 22.95 28.47 29.13 24.09 26.72
Valley[44] 8 LLaMA-13B 28.70 25.00 21.86 30.51 26.38 27.73 27.02

### v2

@@ I Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [yin2023survey, team2023gemini, GPT4v, zhang2023llama, huang2023language, zhu2023minigpt, vlmsurvey] has demonstrated remarkable comprehension and generalization capacities, opening new possibilities for achieving the ultimate goal of artificial general intelligence (AGI) [y2023artificial, morris2023levels], which aims to match or surpass human performance in most tasks. By plugging efficient visual encoders into pretrained Large Language Models (LLMs) [chowdhery2023palm, touvron2023llama, chiang2023vicuna, kenton2019bert] and learning alignments between vision and language [x2vlm], MLLMs have excelled in various multimodal tasks such as image captioning [chen2015microsoft, li2024multimodal], visual question answering [li2024multimodal, antol2015vqa, li2024mvbench, fu2024video, li2024seed, he2024mmworld, zhang2024q], mathematical reasoning [lu2023mathvista, zhang2024mathverse], crossing-modality grounding [peng2023kosmos], etc. However, achieving AGI requires more than just advanced comprehension and reasoning. A crucial milestone is attaining human-level task planning capabilities, which involve making informed decisions in complex environments. This capability is essential for developing a versatile intelligent assistant that can assist humans in tackling a wide array of real-world challenges in daily life.
@@
While the comprehension capabilities of MLLMs have been extensively evaluated in previous benchmarks [liu2023mmbench, li2023seed, li2024seed, li2024seedplus, unk-vqa], the evaluation of the planning abilities of current MLLMs in various scenarios remains underexplored. A comprehensive benchmark specifically designed to assess the planning capabilities of MLLMs across a wide range of real-world scenarios is highly demanded to uncover the potential of MLLMs in serving as versatile assistants in the real world. Previous egocentric video question answering (QA) benchmarks [fan2019egovqa, jia2022egotaskqa] also evaluate model performance in everyday life, but they primarily assess comprehension rather than planning, where a model answers questions based on the spatial and temporal understanding of the entire video. Most relevant to addressing this issue is EgoPlan-Bench [chen2023egoplan], which evaluates the planning abilities of MLLMs from an egocentric perspective. However, it is constrained to single kitchen scenarios, lacking a comprehensive evaluation across a variety of real-world contexts.
@@
In this paper, we introduce EgoPlan-Bench2, a benchmark designed to rigorously assess the planning capabilities of MLLMs across a broad range of daily scenarios. EgoPlan-Bench2 is founded on three principal tenets: a) Rich and diverse real-world scenarios. It includes 1,321 high-quality multiple-choice QA pairs sourced from 1,113 videos, covering 4 major life domains: Work, Daily life, Hobbies and Recreation. These domains are further subdivided into 24 detailed scenarios, ranging from everyday household tasks to specialized activities such as laboratory work, blacksmith and mechanical repairs as shown in Fig. 1. In addition to scenario settings, EgoPlan-Bench2 features 284 distinct verbs in task goals and 434 in candidate options, along with 742 and 1,113 unique objects respectively. The duration of task progress videos varies from a few seconds to five minutes. The rich and diverse scene setups, coupled with a variety of actions, objects and video lengths, ensure a comprehensive evaluation of MLLMs’ planning capabilities across various contexts. b) Egocentric perspective. We choose Ego4D [grauman2022ego4d] as the video source, because it provides a vast array of first-person perspective videos that capture realistic human interactions with objects and environments. This egocentric perspective is crucial for evaluating planning capabilities in a manner that closely mirrors real-world human experiences, offering a more authentic assessment of how MLLMs can assist in everyday tasks. c) Planning capability evaluation. EgoPlan-Bench2 is specifically aimed at evaluating MLLMs’ planning abilities, where a model must track long-term task progress, comprehend the current state of the environment, and leverage both general and domain-specific knowledge to plan the next action, in order to correctly answer the questions as shown in Fig. 1.
@@ II-A Advancements in Multimodal Large Language Models
Building upon the impressive achievements of LLMs, MLLMs have also experienced a revolutionary transformation. MLLMs typically consists of an image encoder (e.g., CLIP [radford2021learning]) to extract visual information, a language model (e.g., LLaMA [touvron2023llama], Vicuna [chiang2023vicuna]) to decode multimodal or text sequence and a trainable align module (e.g., Q-Former [li2023blip], gated cross-attention layer [alayrac2022flamingo]) to integrate visual features into the language embedding space. Groundbreaking models like BLIP-2 [li2023blip], LLaVA [liu2024visual], Flamingo [alayrac2022flamingo] and PaLM-E [driess2023palm] have made early attempt to integrate LLMs into vision-language pre-training and have demonstrated remarkable multimodal understanding and reasoning capabilities. Recent research interest has increasingly shifted towards multimodal understanding and generation that incorporates videos as visual signals [li2023videochat, maaz2023video, ge2023planting, ge2023making, ge2024seed]. These methods (e.g., VideoChat [li2023videochat], VideoChatGPT [maaz2023video], Valley [luo2023valley]) try to enhance MLLMs’ instruction-following capabilities by generating video instruction-tuning data. Video-LLaMA [zhang2023video] encodes individual frames through a ViT [dosovitskiy2020image] and an image Q-Former and then apply temporal process through a video Q-Former. VideoChat2 [li2024mvbench] encodes video frames through a video transformer, and Q-former is employed to compress video tokens. Expect these models mentioned above, a considerable number of video-based MLLMs [chen2024sharegpt4video, zhang2024llavanextvideo, zhang2024longva, cheng2024videollama] have been proposed, demonstrating notable generalization and reasoning abilities across a wide range of tasks.
@@ II-B Benchmarking Multimodal Large Language Models
To guide the potential future development of MLLMs, numerous benchmarks have been proposed to assess model performance across various aspects and tasks. Alongside advancements in comprehensive benchmarks for image MLLMs [yu2023mm, liu2023mmbench, xu2023lvlm, ye2023mplug, yin2024lamm, yue2024mmmu, ying2024mmt], significant efforts have been made to creating benchmarks for video MLLMs [patraucean2024perception, maaz2023video, fu2024video, li2024mvbench, he2024mmworld]. For example, Video-MME [fu2024video] is build upon videos collected from YouTube, aiming to evaluate models’ capacities in 12 task types involving temporal perception, spatial perception, OCR, object recognition, etc. MVBench [li2024mvbench] focuses on temporally-sensitive videos and conducts comprehensive evaluations of MLLMs’ temporal understanding. MMWorld [he2024mmworld] is a benchmark characterized by its multi-discipline nature, evaluating models on multi-faceted tasks such as explanation and counterfactual thinking. MLVU [zhou2024mlvu] is a long video understanding benchmark which is built upon various video genres, including movies, cartoons, game videos, etc. The diagnostic benchmark dataset, WorldQA [zhang2024worldqa], challenges machines to answer questions about a video by employing multimodal data (auditory and visual) and world knowledge. Despite the advancement, videos in these benchmarks are not recorded from a first-person perspective and fail to simulate the realistic visual input that a model would receive during planning tasks. In addition, these benchmarks mainly focus on MLLMs’ comprehension capabilities instead of planning capabilities.
@@ II-C Egocentric Video Datasets
Numerous egocentric datasets containing daily life activity have been developed over the past years, including Ego4D [grauman2022ego4d], Epic-Kitchens [epickitchen], UT Ego [lee2012discovering, su2016detecting], Activities of Daily Living (ADL) [pirsiavash2012detecting], Disney dataset [fathi2012social], Charades-Ego [sigurdsson2018charades], etc. Among these egocentric datasets, we select Ego4D as the video source for EgoPlan-Bench2 due to its massive scale, encompassing hundreds of indoor and outdoor environments, a much wider demographic and an exceptionally rich variety of tasks and scene types. There are also some existing egocentric QA benchmarks like EgoThink [cheng2023can], EgoVQA [fan2019egovqa] and Egotaskqa [jia2022egotaskqa], but they mainly focus on evaluating model’s reasoning and comprehension abilities of activities, human-object interactions and environments instead of task planning capacity. Most relevant to our work are EgoPlan-Bench [chen2023egoplan] and VidEgoThink [cheng2024videgothink]. EgoPlan-Bench is also a planning benchmark that utilizes kitchen-related videos from Epic-Kitchens [epickitchen] and Ego4D datasets, but it focuses exclusively on cooking scenario and lacks evaluation across a wider variety of real task types. VidEgoThink evaluates the capabilities for different functions of MLLMs in Embodied AI from four dimensions: video QA, hierarchy planning, visual grounding and reward modeling. In the most relevant dimension of hierarchy planning, VidEgoThink includes only 9 scenes sourced from 76 videos, suffering from limited task types and scenario diversity.
@@ III-A Egocentric Video Source
Ego4D has densely timestamped annotations indicating specific actions and their occurrence time, represented by short sentences in the format of “#C C does something”. In order to remove invalid annotations and reduce noise, we adopt the criteria following the EgoVLP framework [lin2022egocentric]:
@@
After narration filtering, we convert raw narrations to verb-object phrases like “close washing machine” with GPT-4 to better represent actions and uniform format. The raw action timestamps in Ego4D indicate when actions occur but not their durations. Following EgoVLP [lin2022egocentric], we calculate start and end timestamps for each action. For an action with occurrence timestamp $t_{i}$, the start and end timestamps can be calculated as:
@@ III-B3 Stage III: Model and Human Verification
In the final verification phase of EgoPlan-Bench2, we exclude QA pairs that can be correctly answered using only textual information, ensuring the benchmark effectively assesses the models’ capabilities to process multimodal inputs. GPT-4 is employed with the CircularEval Strategy [liu2023mmbench] to minimize the impact of random guessing.
@@ III-C EgoPlan-Bench2 Statistics
EgoPlan-Bench2 is highly diverse in terms of scenarios, human-object interactions, and video durations, comprising 1,321 high-quality multiple-choice QA pairs derived from 1,113 videos. Firstly, it has an extensive range of scenario settings across 4 major life domains: Work (404 QA pairs), Daily Life (439 QA pairs), Recreation (183 QA pairs) and Hobbies (295 QA pairs). A more detailed breakdown of the 24 fine-grained scenarios and the corresponding number of QA pairs is provided in Fig. 4. Second, EgoPlan-Bench2 incorporates comprehensive human-object interaction data, manifested in both actions and task goals. We use NLTK [bird2009natural] to analyze the verbs and nouns in the task goals and candidate options of the QA pairs, identifying 284 distinct verbs and 742 unique objects in the task goals, and 434 verbs and 1,113 objects in the candidate options. The word clouds of these terms are displayed in Fig. 5. Third, varied video durations in EgoPlan-Bench2 range from a few seconds to five minutes. We consider 30 seconds as the threshold, classifying videos of 30 seconds or less as short videos (690 QA pairs) and those exceeding this duration as long videos (631 QA pairs).
@@ IV Experiments
Yi-VL[young2024yi] 8 Yi-6B 24.37 21.29 26.23 23.39 23.19 23.77 23.47
MultiModal-GPT[gong2023multimodal] 8 LLaMA-7B 26.42 23.27 23.50 26.10 23.48 26.62 24.98
LLaVA1.5[liu2024improved] 6 LLaMA-7B 29.61 21.04 27.32 24.07 23.77 27.26 25.44
InternVL-1.5[chen2024internvl, chen2024far] 8 InternLM2-Chat-1.8B 28.02 24.75 23.50 24.41 23.91 27.42 25.59
mPLUG-Owl-2[ye2024mplug] 8 LLaMA2-7B 27.79 24.75 24.04 25.42 25.36 26.31 25.81
BLIP-2[li2023blip] 8 Flan-T5-XL 24.37 23.51 30.05 30.17 24.64 27.89 26.19
InstructBLIP[dai2023instructblip] 8 Flan-T5-XL 27.33 23.02 26.23 29.49 25.51 27.26 26.34
InstructBLIP Vicuna[dai2023instructblip] 8 Vicuna-7B 27.56 24.26 28.42 28.14 25.65 28.05 26.80
DeepSeek-VL[lu2024deepseek] 6 DeepSeek-LLM-7B 32.12 24.75 26.23 29.83 28.55 28.53 28.54
Qwen-VL-Chat[bai2023qwen] 8 Qwen-7B 32.57 27.23 27.87 28.47 30.00 28.68 29.37
InternVL-2[chen2024internvl, chen2024far] 8 InternLM2.5-Chat-7B 37.81 23.76 31.69 28.14 31.01 29.95 30.51
@@
Video-LLaMA2[cheng2024videollama] 8 Mistral-v0.2-Instruct-7B 24.15 23.02 19.13 23.73 23.19 22.82 23.01
LLaVA-NeXT-Video[zhang2024llavanextvideo] 16 Vicuna1.5-7B 26.42 19.55 24.59 23.05 22.61 24.09 23.32
Video-ChatGPT[maaz2023video] 100 LLaMA-7B 24.15 22.77 24.59 24.07 23.33 24.25 23.77
Video-LLaVA[lin2023video] 8 Vicuna1.5-7B 27.11 22.52 27.87 24.75 25.22 25.36 25.28
ShareGPT4Video[chen2024sharegpt4video] 16 LLaMA3-Instruct-8B 25.51 23.02 26.78 27.46 25.07 25.67 25.36
VILA[lin2024vila] 6 LLaMA3-8B 28.70 20.05 30.05 25.08 23.77 27.26 25.44
LongVA[zhang2024longva] 32 Qwen2-Instruct-7B 27.11 23.27 26.78 29.49 27.25 25.52 26.42
VideoChat2[li2024mvbench] 16 Mistral-v0.2-Instruct-7B 28.93 24.75 22.95 28.47 29.13 24.09 26.72
Valley[luo2023valley] 8 LLaMA-13B 28.70 25.00 21.86 30.51 26.38 27.73 27.02